# Global magnetohydrodynamic simulations of the inner regions of protoplanetary discs. I. Zero-net flux regime

### v1

@@ 6.3 Formation mechanism
We now speculate on the formation mechanism of the large-scale magnetic field structures in the dead zone, as examined in the preceding parts of this section. We distinguish between different proposed formation pathways for large-scale magnetic field structures, dividing our analysis into the development of the poloidal and toroidal components.
@@
Using Fig. 19 as a reference, we propose the following sequence. Turbulent diffusion in the active zone transports ordered poloidal magnetic fields outward across the dead–active zone interface. Beyond this transition, non-ideal MHD diffusion controls the transport of these fields. The outward transport is stronger in the disc surface layers because the diffusion coefficients at the disc surface (dominated by $\eta_{A}$) are greater than at the midplane (dominated by $\eta_{O}$): $\eta_{A}(z=4H)/\eta_{O}(z=0)\sim\text{R}_{m}(z=0)/\beta(z=4H)>1$. Furthermore, the field lines in the surface layers are more bent, leading to shorter local length scales, and faster diffusive speeds. Together, these effects produce a bow-like structure, with field lines curving back toward the midplane near the dead–active zone interface (left panel). The inversion point subsequently develops at the midplane, likely due to reconnection (middle panel). The resulting complex field structure then expands and gradually diffuses outward (see Section 6.2.1), ultimately forming the concentric large-scale poloidal field loops (right panel).
Second, we examine the development of toroidal magnetic fields in the dead zone. Meridional plots of $R\langle{B_{\phi}}\rangle$ in Fig. 5 and Fig. 19, show that $B_{\phi}$ gradually emerges within the dead zone, and that its generation is broadly anti-correlated with the sign of $B_{r}$. These trends suggest that the toroidal field is primarily generated by the $\Omega$-effect acting on $B_{r}$, rather than through the outward diffusion of toroidal flux from the active zone. The characteristic timescale for winding by the $\Omega$-effect ($\sim\Omega_{K}^{-1}$) is shorter than the Ohmic diffusion timescale, as the magnetic Reynolds number is relatively large ($\text{R}_{m}\gg 1$), with the initial $\text{R}_{m}$ profile shown in Fig. 1. Consequently, the generation of $B_{\phi}$ can be roughly approximated by $B_{\phi}(t)\sim-B_{r}\Omega t$. In the outer midplane region of the disc, where magnetic diffusion is weakest and beyond the inversion point, the toroidal field develops over the interval $t_{\text{in}}\in[200,700]$ (see Fig. 3). During this period, $B_{r}$ remains approximately steady at $\sim 5\times 10^{-5}$, leading to an expected $B_{\phi}$ amplitude of $\sim 10^{-3}$ after $\sim 25$ local orbits, consistent with values measured from the simulation. For toroidal fields that grow over longer timescales (see the dead zone surface layers in Fig. 19), amplification is ultimately mediated by ambipolar diffusion, which scales with $B^{2}$.
@@ 7 Conclusion
We performed five three-dimensional global ZNF non-ideal MHD simulations, to investigate the dynamics around the dead–active zone interface. These simulations include physically motivated prescriptions for Ohmic and ambipolar diffusion, as well as a disc–corona temperature transition. Our analysis focuses on three key aspects: the structure of the inner disc, the formation and morphology of large-scale hydrodynamic features at the dead–active zone interface, and the evolution of large-scale coherent magnetic fields in the dead zone. Below, we summarise our main results:
@@
    The ZNF MRI dynamo generates coherent large-scale poloidal field loops that expand to fill the entire MRI active region, supporting the phenomenology proposed by Jacquemin-Ide et al. (2024). Our results also highlight the potential role of the dead zone in mediating the outward propagation of this generated field, a regime not explored in previous studies.
@@
    These large-scale poloidal loops are vertically bounded by the disc–corona temperature transition in the MRI active region, leading to the accumulation of strong, tightly wound magnetic fields in the surface layers. This modifies the accretion profile in the active region, producing a significant component of surface layer accretion. We also demonstrated that this behaviour persists in fully MRI active discs, suggesting that the vertical buoyancy configuration imposed by a corona transition plays a more general role in shaping the disc dynamics. Finally, we speculate that incorporating a sufficiently heated corona could significantly influence the elevated disc structures seen in global vertical-net flux MRI simulations (Zhu & Stone, 2018; Jacquemin-Ide et al., 2021).
@@
    The formation of an axisymmetric pressure maximum that extends through through multiple disc scale heights, is driven by a localised density enhancement, resulting from sustained outward mass flux at the outer edge of the MRI active region.
@@
    Rossby wave-induced vortices consistently form at the dead–active zone interface in the ZNF regime. They encroach upon the active region and are only weakly coherent due to turbulent disruption at the interface. These vortices eventually dissipate, primarily due to numerical dissipation. The robustness of the formation mechanism for both large-scale HD structures, in the presence of an initial vertical-net flux, will be investigated in the second paper of this series.
@@
    The ZNF toroidal magnetic field, initially confined to the active region, develops a complex magnetic morphology that gradually permeates the entire dead zone. Notably, a coherent large-scale vertical magnetic field emerges, threading the inner dead zone of the disc. Although the formation mechanism and robustness of this configuration remain uncertain, this represents the first instance of a self-consistently generated magnetic structure capable of launching a weak magnetic pressure wind from the dead zone in protoplanetary disc simulations.
Global simulations provide valuable insight, but they come with important caveats. They isolate a section of the disc from its broader environment, evolve it from idealised initial conditions, and typically span only a short timescale relative to the disc’s secular evolution. Moreover, their ability to systematically explore the parameter space – such as boundary conditions, domain size, temperature prescriptions, and initial magnetic configurations – is inherently limited. These issues are particularly relevant in the inner disc, where our thermodynamically simple results must be considered in the context of the observed variability, and the connection to the complex disc–star boundary (e.g. Zhu, 2025; Takasao et al., 2025).
Whilst we have proposed mechanisms for the development of large-scale magnetic field structures in our simulations – including surface-layer accretion near the disc–corona transition in the active zone, and the permeation of the dead zone by magnetic fields – both of these intriguing phenomena merit further investigation in more controlled settings.
In addition to these outstanding questions, our results suggest several promising directions for future work within both the ZNF and vertical-net flux regimes. Key extensions include incorporating the Hall effect, due to its known dynamical influence in the dead zone (Wurster, 2021); implementing a dust–gas feedback module to explore dust aggregation in the large-scale HD structures; and introducing a mobile dead–active zone interface (e.g. Faure et al., 2015), to probe the interpenetration of dynamics and thermodynamics in this region.
However, our immediate next step is to extend the present study to the vertical-net flux regime in the second paper within this series, similar to that recently explored by Iwasaki et al. (2024). There, we focus on global magnetic flux transport, strong variability at the dead–active zone interface, and the interplay between laminar and turbulent flows in the inner disc – particularly in relation to outflows, accretion, and the formation of large-scale hydrodynamic structures.

### v3

@@ 6.3 Formation mechanism
We now speculate on the formation mechanism of the large-scale magnetic field structures in the dead zone, as examined in the preceding parts of this section. We distinguish between different proposed formation pathways, dividing our analysis into the development of the poloidal and toroidal components separately.
@@
Using Fig. 19 as a reference, we propose the following sequence. Turbulent diffusion in the active zone transports ordered poloidal magnetic fields outward across the dead–active zone interface. Beyond this transition, non-ideal MHD diffusion controls the transport of these fields. The outward transport is stronger in the disc surface layers because the diffusion coefficients at the disc surface (dominated by $\eta_{\text{A}}$) are greater than at the midplane (dominated by $\eta_{\text{O}}$): $\eta_{\text{A}}(z=4H)/\eta_{\text{O}}(z=0)\sim\text{R}_{\text{m}}(z=0)/\beta(z=4H)>1$. Furthermore, the field lines in the surface layers are more bent, leading to shorter local length scales, and faster diffusive speeds. Together, these effects produce a bow-like structure, with field lines curving back towards the midplane near the dead–active zone interface (left panel). The inversion point subsequently develops at the midplane, likely due to reconnection (middle panel). The resulting complex field structure then expands and gradually diffuses outward (see Section 6.2.1), ultimately forming the concentric large-scale poloidal field loops (right panel).
Second, we examine the development of toroidal magnetic fields in the dead zone. The meridional plots of $R\langle{B_{\phi}}\rangle_{\phi,t}$ in Fig. 5 and Fig. 19 show that $B_{\phi}$ gradually emerges within the dead zone, and that its generation is broadly anti-correlated with the sign of $B_{r}$. These trends suggest that the toroidal field is primarily generated by the $\Omega$-effect acting on $B_{r}$, rather than through the outward diffusion of toroidal flux from the active zone. The characteristic timescale for winding by the $\Omega$-effect ($\sim\!\Omega_{\text{K}}^{-1}$) is shorter than the Ohmic-diffusion timescale, as the magnetic Reynolds number is relatively large ($\text{R}_{\text{m}}\gg 1$), with the initial $\text{R}_{\text{m}}$ profile shown in Fig. 1. Consequently, the generation of $B_{\phi}$ can be roughly approximated by $B_{\phi}(t)\sim-B_{r}\Omega t$. In the outer midplane region of the disc, where magnetic diffusion is weakest and beyond the inversion point, the toroidal field develops over the interval $t_{\text{in}}\in[200,700]$ (see Fig. 3). During this period, $B_{r}$ remains approximately steady at $\sim\!5\times 10^{-5}$, leading to an expected $B_{\phi}$ amplitude of $\sim\!10^{-3}$ after $\sim\!25$ local orbits, consistent with the values measured from the simulation. For toroidal fields that grow over longer timescales (see the dead-zone surface layers in Fig. 19), amplification is ultimately mediated by ambipolar diffusion, which scales with $B^{2}$.
@@ 7 Conclusion
We performed five three-dimensional global ZNF non-ideal MHD simulations, to investigate the dynamics around the dead–active zone interface. These simulations used physically motivated prescriptions for Ohmic and ambipolar diffusion, as well as a disc–corona temperature transition. Our analysis focused on three key aspects: the structure of the inner disc, the formation and morphology of large-scale hydrodynamic features at the dead–active zone interface, and the evolution of large-scale coherent magnetic fields in the dead zone. Below, we summarise the main results:
@@
    The ZNF MRI dynamo generates coherent, large-scale poloidal field loops that expand to fill the entire MRI-active region, supporting the phenomenology proposed by jacquemin-ide_magnetorotational_2024. Our results also highlight the potential role of the dead zone in mediating the outward propagation of this generated field, a regime not explored in previous studies.
@@
    These large-scale poloidal loops are vertically bounded by the disc–corona temperature transition in the MRI-active region, leading to the accumulation of strong, tightly wound magnetic fields in the surface layers. This modifies the accretion profile in the active region, producing a significant component of surface-layer accretion. We also demonstrate that this behaviour persists in fully MRI-active discs, suggesting that the vertical buoyancy configuration imposed by a corona transition plays a broader role in shaping disc dynamics. Finally, we speculate that incorporating a sufficiently heated corona could significantly influence the elevated disc structures seen in global vertical-net flux MRI simulations (zhu_global_2018; jacquemin-ide_magnetic_2021).
@@
    The formation of an axisymmetric pressure maximum that extends through multiple disc scale heights is driven by a localised density enhancement, resulting from sustained outward mass flux at the outer edge of the MRI-active region.
@@
    Rossby-wave-induced vortices consistently form at the dead–active zone interface in the ZNF regime. They encroach upon the active region and are only weakly coherent due to turbulent disruption at the interface. These vortices eventually decay, primarily through numerical dissipation. The robustness of the formation mechanism for both large-scale HD structures in the presence of an initial vertical-net flux will be investigated in the second paper of this series.
@@
    The ZNF toroidal magnetic field, initially confined to the active region, develops a complex morphology that gradually permeates the entire dead zone. Notably, a coherent large-scale vertical magnetic field emerges, threading the inner dead zone of the disc. Although the formation mechanism and robustness of this configuration remain uncertain, this represents the first instance of a self-consistently generated magnetic structure capable of launching a weak magnetic-pressure wind from the dead zone in protoplanetary disc simulations.
Global simulations provide valuable insight, but come with important caveats. They isolate a section of the disc from its broader environment, evolve it from idealised initial conditions, and typically span only a short timescale relative to the disc’s secular evolution. Moreover, their ability to systematically explore the parameter space – such as boundary conditions, domain size, temperature prescriptions, and initial magnetic configurations – is inherently limited. These issues are particularly relevant in the inner disc, where our thermodynamically simple results must be considered in the context of the observed variability, and the connection to the complex disc–star boundary (e.g. zhu_global_2025; takasao_connecting_2025).
Whilst we have proposed mechanisms for the development of large-scale magnetic field structures in our simulations – including surface-layer accretion near the disc–corona transition in the active zone, and the permeation of the dead zone by magnetic fields – both of these intriguing phenomena merit further investigation in more controlled settings. In particular, the generation of large-scale poloidal field loops within the active region at smaller disc aspect ratios, and the influence of a hot corona on the MRI-driven dynamo, both warrant further study.
In addition to these outstanding questions, our results suggest several promising directions for future work within both the ZNF and vertical-net flux regimes. Key extensions include incorporating the Hall effect, due to its known dynamical influence in the dead zone (wurster_we_2021); implementing a dust–gas feedback module to explore dust aggregation in the large-scale HD structures; and introducing a mobile dead–active zone interface (e.g. faure_vortex_2015), to probe the interpenetration of dynamics and thermodynamics in this region.
However, our immediate next step is to extend the present study to the vertical-net flux regime in the second paper within this series, similar to that recently explored by iwasaki_dynamics_2024. There, we focus on global magnetic flux transport, strong variability at the dead–active zone interface, and the interplay between laminar and turbulent flows in the inner disc – particularly in relation to outflows, accretion, and the formation of large-scale hydrodynamic structures.
@@ Acknowledgements
The authors would like to thank the anonymous referee for useful comments that helped improve the manuscript.